OpenStudy (melstutes):

2x + Y=8 Y=x-7 solve by substitution

OpenStudy (freckles):

replace the Y in the first equation with x-7

OpenStudy (freckles):

2x+Y=8 so if Y=x-7 then 2x+(x-7)=8 <--solve for x

OpenStudy (anonymous):

If you know what y equals, then you can put what it equals into the equation in place of every y you see. in this case, if y=x-7, then the equation would come out to 2x+x-7=8, therefore, 3x=1, x = 1/3. It is called substitution because you substitute what you know for what you want to know.

OpenStudy (dtan5457):

Yeah...2x+(x-7)=8 just plug the given y to x..... you end up with 3x-7=8, x=5